# Surfaces and other Peano Continua with no Generic Chains

### v1

@@ 2. Peano Continua and Chains
A continuum is a compact connected metrizable space. A Peano continuum is a locally connected continuum. An arc is a homeomorphic image of the interval $[0,1]$; a continuum theoretical graph is a set composed of finitely many arcs that intersect each other at most in their endpoints; a continuum theoretical tree is a connected continuum theoretical graph that contains no simple closed curve, meaning no homeomorphic copy of $S^{1}$.
@@
Let $\mathcal{V}$ be another family of open subsets of $X$ and, for $V\in\mathcal{V}$, let
@@
Note that if $\mathcal{W}$ is a family of connected open sets, then $\operatorname{Core}(V,\mathcal{W})=\mathcal{W}(V)\setminus\operatorname{Star}(\operatorname{Bd}(V),\mathcal{W})$.
An open set $U$ is regular if $U=\operatorname{Int}(\operatorname{Cl}(U))$. If $U,V\subseteq X$ are regular open sets, let $U\vee V$ be the regular open set $\operatorname{Int}(\operatorname{Cl}(U)\cup\operatorname{Cl}(V))$ and, similarly, let $\bigvee\mathcal{U}\coloneqq\operatorname{Int}(\bigcup_{U\in\mathcal{U}}\operatorname{Cl}(U))$, for any family $\mathcal{U}$ of regular open sets.
@@ 2.1. The Space of Maximal Connected Chains
Given a regular topological space $X$, let $K(X)$ denote the set of compact subspaces of $X$. The Vietoris topology on $K(X)$ is generated by sets obtained from nonempty regular open subsets $U_{0},\dots,U_{n-1}$ of $X$ as
@@
If $X$ is compact, the Vietoris topology on $K(X)$ is also compact and the subspace $C(X)\subseteq K(X)$ of subcontinua, i.e. compact connected sets, is closed (hence compact itself). If $X$ is locally connected, a basis for this topology on $C(X)$ is given by sets of the form $O(U_{0},\dots,U_{n-1})$, where $U_{i}$ are open and connected and such that $\bigvee_{i<n}U_{i}$ is also connected.
A collection $\mathcal{C}$ of compact subsets of $X$ is called a chain (a nest in [illanes1999hyperspaces]) if it is linearly ordered by inclusion, that is, if for every $C_{1},C_{2}\in\mathcal{C}$, either $C_{1}\subseteq C_{2}$ or $C_{2}\subseteq C_{1}$. A chain $\mathcal{C}\subseteq K(X)$ is maximal if it is not properly contained in any other chain and it is connected if it consists of connected sets.
@@
which is a compact subset of $C(C(X))$ (see [uspenskij2000universal]*§2). The map $\Phi(X)\to X$ that that to each chain $\mathcal{C}$ associates the unique element of the singleton $\bigcap\mathcal{C}$ – the so called root of $\mathcal{C}$ – is continuous.
@@
A $\pi$-basis $\mathcal{B}$ of a topological space $X$ is a collection of open subsets of $X$ such that for every open $U\subseteq X$ there exists $B\in\mathcal{B}$ with $B\subseteq U$. Given a $\pi$-basis $\mathcal{B}$ of $X$, one can construct a $\pi$-basis of $\Phi(X)$ by considering sets of the form described below as $\mathcal{B}_{0},\dots,\mathcal{B}_{n-1}$ vary among all finite subsets of $\mathcal{B}$:
$\Phi(X)\cap\bigcap_{i<n}O\left(O(\mathcal{B}_{0}),\dots,O(\mathcal{B}_{i})\right)=\{\mathcal{C}\in\Phi(X):\mathcal{C}\cap O(\mathcal{B}_{i})\neq\emptyset\text{ for every }i<n\}.$
@@ 3. Walks on Graphs
Throughout this paper, the word graph refers to finite connected reflexive graphs.
@@
A function $f\colon(G,E)\to(H,E^{\prime})$ is a monotone epimorphism (also called connected epimorphism in [PS22]) if it is a homomorphism of graphs such that the preimage of each nonempty connected subset of $H$ is nonempty and connected in $G$. Equivalently, if it is a homomorphism of graphs which is surjective on edges and such that the preimage of each vertex of $H$ is nonempty and connected (see [PS22]*Lemma 1.1). In particular, monotone epimorphisms are onto.
@@
A walk is circular if $w(-1)=w(0)$. We modify the definitions of path/spaced for circular walks accordingly. We also modify the definition of length, so that $\operatorname{lh}(w)=\ell$ if $w=\langle w(0),\dots,w(\ell-1),w(0)\rangle$. This is off-by-one with respect to its length as a walk.
Given two walks $w,v$ on $(G,E)$, we say that $v$ is confined in $w$ if
$\{v(j):j<\operatorname{lh}(v)\}\subseteq\{w(i):i<\operatorname{lh}(w)\}$
and that it is confined in some $A\subseteq G$ if $\{v(j):j<\operatorname{lh}(v)\}\subseteq A$. We say that $w$ refines $v$, in symbols $w\sqsubseteq v$, if:
@@
Finally, we will say that a walk is a lasso if it can be written as $v{}^{\frown}c$, where $v$ is a walk and $c$ is a circular walk; it is a (spaced) lasso path if moreover $v$ is a (spaced) path and $c$ is a (spaced) circular path disjoint from $v$. When we say that a lasso $v{}^{\frown}c$ monotonically refines another $v^{\prime}{}^{\frown}c^{\prime}$, we additionally require that $v$ is confined to $v^{\prime}$ and $c$ to $c^{\prime}$.
@@ 3.1. The Winding Number of a Walk
Let $(G,E)$ be a graph and $C$ be a circular path on $G$ of length $\ell$. For each ordered pair of vertices $(x,y)$ in $G$, we define its weight relative to $C$ to be:
@@
Notice that $\operatorname{weight}_{C}$ is well defined since $C$ is a path. For any walk $w$ on $G$, we define its winding number around $C$ as
@@
We devote the end of the section to a few easy lemmas on walks and their winding numbers which will be applied later in the paper. The impatient reader is advised to skip them and return to consult them during the proof of Theorem 6.4. The keen reader, on the other hand, is encouraged to compare Lemma 3.3 with [GTZ]*Lemma 2.12 and Lemma 3.4 with [GTZ]*Lemma 2.10. For ease of reference, the graphs are called $\mathcal{S}$ and $\mathcal{W}$ rather than $(G,E)$ or $(H,E)$, as elsewhere in this section.
@@
$\operatorname{wd}_{C}(w)$ $=\sum_{j<\operatorname{lh}(w)-1}\operatorname{weight}_{C}(w(f(j)),w(f(j+1)))$
@@
Let $\mathcal{S}$ be a graph and $C$ be a spaced circular path on $\mathcal{S}$ of length $\ell\geq 4$. Let $z,z^{\prime}$ be walks of the same length $k$ on $\mathcal{S}$, which are confined to $C$ and such that, for each $i<k$, $d_{\mathcal{S}}(z(i),z^{\prime}(i))\leq 1$, and also, for $i<k-1$, $d_{\mathcal{S}}(z^{\prime}(i),z(i+1))\leq 1$. Then $\lvert\operatorname{wd}_{C}(z)-\operatorname{wd}_{C}(z^{\prime})\rvert\leq 2$.
@@
(3.1) $\operatorname{wd}_{C}(\langle z^{\prime}(0),\dots,z^{\prime}(i)\rangle)=\operatorname{weight}_{C}(z^{\prime}(0),z(0))+\operatorname{wd}_{C}(\langle z(0),\dots,z(i)\rangle)+\operatorname{weight}_{C}(z(i),z^{\prime}(i)).$
@@ 4. Brick Partitions
A metric space $X$ is uniformly locally connected if for every $\varepsilon>0$ there exists $\delta>0$ such that for all $x,y\in X$, $d(x,y)<\delta$ implies that there is a connected set $C\subseteq X$ with $x,y\in C$ and $\operatorname{diam}(C)<\varepsilon$. When $X$ is a Peano continuum, $C$ can be assumed to be an arc.
@@
A family of open sets $\mathcal{W}$ is said to refine another one $\mathcal{V}$, in symbols $\mathcal{W}\preceq\mathcal{V}$, if every element of $\mathcal{W}$ is contained in (exactly) an element of $\mathcal{V}$. In the case of brick partitions, a refinement $\mathcal{W}\preceq\mathcal{V}$ gives rise to a monotone epimorphism from the nerve graph of $\mathcal{W}$ to that of $\mathcal{V}$, given by mapping $W\mapsto V$ whenever $W\subseteq V$. We denote such map by $\rho^{\mathcal{W}}_{\mathcal{V}}\colon\mathcal{W}\to\mathcal{V}$. Given two brick partitions $\mathcal{W},\mathcal{V}$ and two walks $w,v$ on $\mathcal{W}$ and $\mathcal{V}$ respectively, we write $(\mathcal{W},w)\preceq(\mathcal{V},v)$ whenever $\mathcal{W}\preceq\mathcal{V}$ and $\rho^{\mathcal{W}}_{\mathcal{V}}\cdot w\sqsubseteq v$.
@@
Recall that a subset $K\in X$ is locally separating if there is an open connected set $U\subseteq X$ such that $U\setminus K$ is disconnected, that a space is planar if it admits an embedding in $\mathbb{R}^{2}$ and it is locally non-planar if none of its nonempty open sets are planar. We will use the following facts, which easily follow from the definitions.
@@
    if $K\subseteq X$ is a non-locally separating closed set, then $x$ can be chosen disjoint from $K$.
@@
Let $X$ be a Peano continuum, $\mathcal{W}$ be a brick partition of $X$ and let $W\in\mathcal{W}$.
@@
    If $K\subseteq X$ is a non-locally separating closed set, and $x,y\in\operatorname{Cl}(W)$, there is an arc $\lambda\subseteq W\cup\{x,y\}$ disjoint from $K$ connecting $x,y$.
@@ 5. Walks on Brick Partitions and their Refinements
We now apply the framework developed in Section 3 to walks on nerve graphs of brick partitions of Peano continua. The highlight of this section is Theorem 5.8, where we show that the existence of a generic chain on $X$ implies an off-by-one weak amalgamation principle for walks on (nerve graphs of) brick partitions of $X$. In Section 6 we shall derive our non-existence result in Theorem 1.2 from the failure of such principle. The ingredients of in the proof of Theorem 5.8 are a general criterion for the existence of a comeager orbit due to Rosendal, and two discretization results: Proposition 5.2 for open sets of chains on $X$, and Theorem 5.7 for homeomorphism of $X$.
@@
Let $X$ be a Peano continuum and $\mathcal{W}$ a brick partition on it. Given a walk $w=\langle w(0),\dots,w(n-1)\rangle$ on $\mathcal{W}$, we define the following open subset of $\Phi(X)$:
$\mathbf{O}_{\mathcal{W},w}\coloneqq\left\{c\in\Phi(X):c\cap O\left(w(0),\dots,w(i-1)\right)\not=\emptyset\text{ for every }i<\operatorname{lh}(w)\right\}.$
It is immediate from the definitions that if $w^{\prime},w$ are walks on a brick partition $\mathcal{W}$, then
$\mathbf{O}_{\mathcal{W},w^{\prime}}\subseteq\mathbf{O}_{\mathcal{W},w}\quad\text{ if and only if }\quad w^{\prime}\sqsubseteq w,$
and, more in general, that whenever $\mathcal{W}\preceq\mathcal{V}$ and $w,v$ are walk on $\mathcal{W},\mathcal{V}$, respectively
(5.1) $\mathbf{O}_{\mathcal{W},w}\subseteq\mathbf{O}_{\mathcal{V},v}\quad\text{ if and only if }\quad\rho^{\mathcal{W}}_{\mathcal{V}}\cdot w\sqsubseteq v.$
We also take note that for each walk $w$ on $\mathcal{W}$ there is a reduced walk $\bar{w}$ such that $\mathbf{O}_{\mathcal{W},\bar{w}}=\mathbf{O}_{\mathcal{W},w}$.
By varying $\mathcal{U}$ and taking $u$ to be the empty walk in the next proposition, we see in particular that open sets determined by reduced walks form a basis in $\Phi(X)$.
@@
Let $X$ be a Peano continuum, and fix a brick partition $\mathcal{U}$ and a walk $u$ on $\mathcal{U}$. Let $\mathbf{V}\subseteq\mathbf{O}_{\mathcal{U},u}\subseteq\Phi(X)$ be an nonempty open set. There exist a brick partition and a reduced walk $(\mathcal{V},v)\preceq(\mathcal{U},u)$ such that $\mathbf{O}_{\mathcal{V},v}\subseteq\mathbf{V}$.

### v3

@@ 2. Peano Continua and Chains
A continuum is a compact connected metrizable space. A Peano continuum is a locally connected continuum. An arc is a homeomorphic image of the interval $[0,1]$; a continuum theoretical graph is a space composed of finitely many arcs that intersect each other at most in their endpoints; a continuum theoretical tree is a connected continuum theoretical graph that contains no simple closed curve.
@@
If $V$ is an open set, let
@@
Note that $\operatorname{Core}(V,\mathcal{W})=\mathcal{W}(V)\setminus\operatorname{Star}(\operatorname{Bd}(V),\mathcal{W})$.
An open set $U$ is regular if $U=\operatorname{Int}(\operatorname{Cl}(U))$. If $U,V\subseteq X$ are regular open sets, let $U\vee V$ be the regular open set $\operatorname{Int}(\operatorname{Cl}(U)\cup\operatorname{Cl}(V))$ and, similarly, let $\bigvee\mathcal{U}\coloneqq\operatorname{Int}(\operatorname{Cl}(\bigcup_{U\in\mathcal{U}}U))$, for any family $\mathcal{U}$ of regular open sets.
@@ 2.1. The Space of Maximal Connected Chains
Given a topological space $X$, let $K(X)$ denote the set of compact subspaces of $X$. The Vietoris topology on $K(X)$ is generated by sets of form $\{K\in K(X):K\subseteq U\}$ and $\{K\in K(X):K\cap U\neq\emptyset\}$, for $U$ open in $X$. If $X$ is compact, the Vietoris topology on $K(X)$ is also compact and the subspace $C(X)\subseteq K(X)$ of subcontinua, i.e. compact connected sets, is closed (hence compact itself). If $X$ is moreover metrizable, say by a metric $d$, then the associated Hausdorff metric on $K(X)$, defined by
$d_{H}(K,K^{\prime})\coloneqq\max\left\{\sup_{x\in K}d(x,K^{\prime}),\sup_{x^{\prime}\in K^{\prime}}d(K,x^{\prime})\right\},$
is compatible with the Vietoris topology. Given $U_{0},\dots,U_{n-1}$ regular open sets, we define the open set:
@@
A collection $\mathcal{C}$ of compact subsets of $X$ is called a chain (a nest in [illanes1999hyperspaces]) if it is linearly ordered by inclusion, that is, if for every $C_{1},C_{2}\in\mathcal{C}$, either $C_{1}\subseteq C_{2}$ or $C_{2}\subseteq C_{1}$. A chain $\mathcal{C}\subseteq K(X)$ is maximal if it is not properly contained in any other chain, and it is connected if it consists of connected sets.
@@
which is a compact subset of $C(C(X))$ (see [uspenskij2000universal]*§2). The map $\Phi(X)\to X$ that to each chain $\mathcal{C}$ associates the unique element of the singleton $\bigcap\mathcal{C}$ – the so-called root of $\mathcal{C}$ – is continuous. Given a compatible metric $d$ on $X$, we denote by $d_{\Phi}$ the Hausdorff metric on $\Phi(X)$ relative to $d_{H}$.
@@ 3. Walks on Graphs
Throughout this paper, the word graph refers to finite connected reflexive undirected graphs.
@@
[…]
@@
A walk is a circular path if it has length at least $3$, $w(i)\neq w(j)$ for each $i<j<\operatorname{lh}(w)-1$, and $w(-1)=w(0)$. We modify the definition of spaced for circular paths accordingly. We also define the circular length of a circular path $w$ as $\operatorname{lh}(w)-1$.
Given two walks $w,v$ on $(G,E)$, we say that $v$ is confined in $w$ if $\operatorname{Im}(v)\subseteq\operatorname{Im}(w)$, and that $v$ is confined in some $A\subseteq G$ if $\operatorname{Im}(v)\subseteq A$. We say that $w$ refines $v$, in symbols $w\sqsubseteq v$, if:
@@
Finally, we will say that a reduced uncrossed walk is a lasso if its decomposition is $v{}^{\frown}c$, for a path $v$ and a nonempty circular path $c$.
Monotone maps allow lifting walks, as we see in the following lemma.
Lemma 3.1 (Monotone Lifting).
Let $f\colon(G,E)\to(H,E^{\prime})$ be a monotone epimorphism between graphs, and let $w$ be a reduced walk on $(H,E^{\prime})$. Then there is a reduced walk $v$ on $(G,E)$ such that $f\cdot v$ monotonically refines $w$. Moreover, $v$ can be chosen to be uncrossed or a (spaced/circular/spaced circular) path whenever $w$ is.
Proof.
Denote $A_{i}\coloneqq f^{-1}(w(i))$, for each $i<\operatorname{lh}{w}$. By hypothesis, they are all connected subsets of $G$. Let $P_{1}$ be a path from $A_{1}\cap[A_{0}]_{1}$ to $A_{1}\cap[A_{2}]_{1}$ in $A_{1}$, of minimal length. Since $P_{1}$ is of minimal length, it is spaced, and any $P_{1}(j)$ which is not an endpoint of $P_{1}$ has distance at least $2$ from $A_{0}$ and $A_{2}$.
[…]
(3.1) $d_{G}(P_{i+1}(j),P_{i})\geq 2,\text{ for each }0<j<\operatorname{lh}(P_{i+1})-1.$
Let $x_{0}\in A_{0}$ be adjacent to $P_{1}(0)$ and $x_{-1}\in A_{\operatorname{lh}(w)-1}$ be adjacent to $P_{\operatorname{lh}(w)-2}(-1)$. Then the walk $v\coloneqq\langle x_{0}\rangle{}^{\frown}P_{1}{}^{\frown}\cdots P_{\operatorname{lh}(w)-2}{}^{\frown}\langle x_{-1}\rangle$ is reduced and such that $f\cdot v$ monotonically refines $w$, and it is a path whenever $w$ is.
We prove that $v$ is spaced whenever $w$ is. Let $i_{0},j_{0}<\operatorname{lh}(v)$ with $|i_{0}-j_{0}|>1$ be given, and let $i,j<\operatorname{lh}(w)$ be such that $f(v(i_{0}))=w(i)$ and $f(v(j_{0}))=w(j)$. If $i=j$, then $v(i_{0}),v(j_{0})\in P_{i}$, which is spaced, so $d_{G}(v(i_{0}),v(j_{0}))>1$. If $|i-j|>1$, then $d_{H}(w(i),w(j))>1$, since $w$ is spaced, so also $d_{G}(v(i_{0}),v(j_{0}))>1$. Otherwise, suppose without loss of generality that $j=i+1$. We conclude by (3.1).
If $w$ is a circular path: let $P_{0}$ be a path from $A_{0}\cap[P_{\operatorname{lh}(w)-2}(-1)]_{1}$ to $A_{0}\cap[P_{1}(0)]_{1}$ in $A_{0}$, of minimal length. As above, $P_{0}$ is spaced and any $P_{0}(j)$ which is not an endpoint has distance at least $2$ from $P_{\operatorname{lh}(w)-2}$ and $P_{1}$. The walk $v\coloneqq P_{0}{}^{\frown}P_{1}{}^{\frown}\cdots P_{\operatorname{lh}(w)-2}{}^{\frown}\langle P_{0}(0)\rangle$ is a circular path such that $f\cdot v$ monotonically refines $w$, which is spaced whenever $w$ is.
If $w$ is uncrossed: let $w=w_{0}{}^{\frown}c{}^{\frown}\dots{}^{\frown}c{}^{\frown}s$, with $w_{0}$ a path, $c$ a circular path, and $s$ an initial segment of $c$, be its decomposition. Apply the previous constructions to $w_{0}$ to obtain a path $v_{0}$ in $G$, and to $c$ to obtain a circular path $c^{\prime}$ in $G$. Let $s^{\prime}$ be the shortest initial segment of $c^{\prime}$ such that $f\cdot s^{\prime}$ monotonically refines $s$.
[…]
@@ 3.1. The Winding Number of a Walk
Let $(G,E)$ be a graph and $C$ be a circular path on $G$ of circular length $\ell$. For each ordered pair of vertices $(x,y)$ in $G$, we define its weight relative to $C$ to be:
@@
Notice that $\operatorname{weight}_{C}$ is well-defined since $C$ is a path. For any walk $w$ on $G$, we define its winding number around $C$ as
@@
We devote the end of the section to a few easy lemmas on walks and their winding numbers which will be applied later in the paper. The impatient reader is advised to skip them and return to consult them during the proof of Theorem 6.4. The keen reader, on the other hand, is encouraged to compare Lemma 3.4 with [GTZ]*Lemma 2.12 and Lemma 3.5 with [GTZ]*Lemma 2.10. For ease of reference, the graphs are called $\mathcal{S}$ and $\mathcal{W}$ rather than $(G,E)$ or $(H,E^{\prime})$, as elsewhere in this section.
@@
$\operatorname{wd}_{C}(w)$ $=\sum_{j<\operatorname{lh}(v)-1}\operatorname{weight}_{C}(w(f(j)),w(f(j+1)))$
@@
Let $\mathcal{S}$ be a graph and $C$ be a spaced circular path on $\mathcal{S}$ of circular length $\ell\geq 4$. Let $z,z^{\prime}$ be walks of the same length $k$ on $\mathcal{S}$, which are confined to $C$ and such that, for each $i<k$, $d_{\mathcal{S}}(z(i),z^{\prime}(i))\leq 1$, and also, for $i<k-1$, $d_{\mathcal{S}}(z^{\prime}(i),z(i+1))\leq 1$. Then $\lvert\operatorname{wd}_{C}(z)-\operatorname{wd}_{C}(z^{\prime})\rvert\leq 2$.
@@
(3.2) $\operatorname{wd}_{C}(\langle z^{\prime}(0),\dots,z^{\prime}(i)\rangle)=\operatorname{weight}_{C}(z^{\prime}(0),z(0))+\\ +\operatorname{wd}_{C}(\langle z(0),\dots,z(i)\rangle)+\operatorname{weight}_{C}(z(i),z^{\prime}(i)).$
@@ 4. Brick Partitions
A metric space $X$ is uniformly locally connected if for every $\varepsilon>0$ there exists $\delta>0$ such that for all $x,y\in X$, $d(x,y)<\delta$ implies that there is a connected set $C\subseteq X$ with $x,y\in C$ and $\operatorname{diam}(C)<\varepsilon$. Peano continua are uniformly locally connected, and in a Peano continuum the connected set $C$ can be always assumed to be an arc.
@@
A family of open sets $\mathcal{W}$ is said to refine another one $\mathcal{V}$, in symbols $\mathcal{W}\preceq\mathcal{V}$, if every element of $\mathcal{W}$ is contained in an element of $\mathcal{V}$. In the case of brick partitions, a refinement $\mathcal{W}\preceq\mathcal{V}$ gives rise to a monotone epimorphism from the nerve graph of $\mathcal{W}$ to that of $\mathcal{V}$, given by mapping $W\mapsto V$ whenever $W\subseteq V$. We denote such map by $\rho^{\mathcal{W}}_{\mathcal{V}}\colon\mathcal{W}\to\mathcal{V}$. Given two brick partitions $\mathcal{W},\mathcal{V}$ and two walks $w,v$ on $\mathcal{W}$ and $\mathcal{V}$ respectively, we write $(\mathcal{W},w)\preceq(\mathcal{V},v)$ whenever $\mathcal{W}\preceq\mathcal{V}$ and $\rho^{\mathcal{W}}_{\mathcal{V}}\cdot w\sqsubseteq v$.
@@
Recall that a subset $K\subseteq X$ is locally separating if there is an open connected set $U\subseteq X$ such that $U\setminus K$ is disconnected, that a space is planar if it admits an embedding in $\mathbb{R}^{2}$, and it is locally non-planar if none of its nonempty open sets are planar. We will use the following facts, which easily follow from the definitions.
@@
    if $K\subseteq X$ is a non-locally-separating subset with empty interior, then $x$ can be chosen disjoint from $K$.
@@
Let $X$ be a Peano continuum and $W\subseteq X$ a connected and uniformly locally connected regular open set.
@@
    If $K\subseteq X$ is a non-locally-separating $F_{\sigma}$ set with empty interior, and $x,y\in\operatorname{Cl}(W)$, there is an arc $\lambda\subseteq(W\setminus K)\cup\{x,y\}$ connecting $x,y$.
Proof of 4.3.(4).
By [MOT]*Proposition 2.5, $\operatorname{Bd}(W)$ is non-locally-separating in $\operatorname{Cl}(W)$, thus so is $K\cup\operatorname{Bd}(W)$, see [MOT]*Theorem 2.2. Let $K^{\prime}=(K\cup\operatorname{Bd}(W))\setminus\{x,y\}$. We claim that $K^{\prime}$ is still non-locally-separating in $\operatorname{Cl}(W)$. Indeed, suppose that $U\subseteq\operatorname{Cl}(W)$ is a nonempty open connected set such that $U\setminus K^{\prime}$ is not connected. On the other hand $U\setminus K$ is connected, and this is possible only if $K\cup\operatorname{Bd}(W)$ contains a connected component of $U\setminus K^{\prime}$, which gives a contradiction since both $K$ and $\operatorname{Bd}(W)$ have empty interior. In particular, $K^{\prime}\cap\operatorname{Cl}(W)$ is a non-locally-separating $F_{\sigma}$ subset of $\operatorname{Cl}(W)$. The conclusion then follows by [MR0007095]*II.5.5, which states that all connected, locally connected, $G_{\delta}$ subsets of a complete metric space are arcwise connected. ∎
@@ 5. Walks on Brick Partitions and their Refinements
We now apply the framework developed in Section 3 to walks on nerve graphs of brick partitions of Peano continua. The highlight of this section is Theorem 5.9, where we show that the existence of a generic chain on $X$ implies an off-by-one weak amalgamation principle for walks on (nerve graphs of) brick partitions of $X$. In Section 6 we shall derive our non-existence result in Theorem 1.2 from the failure of such principle.
The ingredients of the proof of Theorem 5.9 are a general criterion for the existence of a comeager orbit due to Rosendal, and two discretization results: Corollary 5.4 for open sets of chains on $X$, and Theorem 5.8 for homeomorphism of $X$. Such discretization process revolves around the next definition, which explains how to associate open subsets of $\Phi(X)$ to walks on brick partitions.
@@
Let $X$ be a Peano continuum and $\mathcal{W}$ a brick partition on it. Given a walk $w$ on $\mathcal{W}$, we define the following open subset of $\Phi(X)$:
$\mathbf{O}_{\mathcal{W},w}\coloneqq\left\{\mathcal{C}\in\Phi(X):\mathcal{C}\cap O\left(w(0),\dots,w(i-1)\right)\not=\emptyset\text{ for every }i<\operatorname{lh}(w)\right\}.$
Notice that $\mathbf{O}_{\mathcal{W},w}$ is nonempty: for each $i<\operatorname{lh}(w)$, choose a point $x_{i}\in w(i)$, and let $\gamma_{i}$ be an arc connecting $x_{i}$ to $x_{i+1}$ which is contained in $w(i)\vee w(i+1)$, which exists by 4.3.(1). Call $K_{0}\coloneqq\{x_{0}\}$ and $K_{i}\coloneqq\bigcup_{j<i}\gamma_{j}$ for each $0<i<\operatorname{lh}(w)$. Then $K_{i}$ is a subcontinuum of $X$ such that $K_{i}\in O\left(w(0),\dots,w(i-1)\right)$, and $K_{i}\subseteq K_{i+1}$. [illanes1999hyperspaces]*Theorem 14.6 then ensures that there is $\mathcal{C}\in\Phi(X)$ such that $K_{i}\in\mathcal{C}$, for each $i<\operatorname{lh}(h)$, so $\mathcal{C}\in\mathbf{O}_{\mathcal{W},w}$.
Also notice that for each walk $w$ on $\mathcal{W}$ there is a reduced walk $\bar{w}$ such that $\mathbf{O}_{\mathcal{W},\bar{w}}=\mathbf{O}_{\mathcal{W},w}$.
It is immediate from the definitions that if $w^{\prime},w$ are walks on a brick partition $\mathcal{W}$,
$\text{if }\quad w^{\prime}\sqsubseteq w\quad\text{ then }\quad\mathbf{O}_{\mathcal{W},w^{\prime}}\subseteq\mathbf{O}_{\mathcal{W},w},$
and, more in general, that whenever $\mathcal{W}\preceq\mathcal{V}$ and $w,v$ are walks on $\mathcal{W},\mathcal{V}$, respectively
(5.1) $\text{if }\quad\rho^{\mathcal{W}}_{\mathcal{V}}\cdot w\sqsubseteq v\quad\text{then }\quad\mathbf{O}_{\mathcal{W},w}\subseteq\mathbf{O}_{\mathcal{V},v}.$
If $\operatorname{Im}(w)\subseteq\operatorname{Im}(w^{\prime})$, we have a strong converse to the above:
(5.2) $\text{if }\quad\mathbf{O}_{\mathcal{W},w^{\prime}}\cap\mathbf{O}_{\mathcal{W},w}\neq\emptyset\quad\text{ then }\quad w^{\prime}\sqsubseteq w.$
Indeed, if $w^{\prime}\not\sqsubseteq w$ but $\operatorname{Im}(w)\subseteq\operatorname{Im}(w^{\prime})$, then there are initial segments $\{w(0),\dots,w(i-1)\},\{w^{\prime}(0),\dots,w^{\prime}(j-1)\}$ of $w,w^{\prime}$, respectively, which are incomparable under inclusion, so no chain can intersect both $O(w(0),\dots,w(i-1))$ and $O(w^{\prime}(0),\dots,w^{\prime}(j-1))$.
In particular, if $w,w^{\prime}$ are two walks with $\operatorname{Im}(w)=\operatorname{Im}(w^{\prime})=\mathcal{W}$, then $w\sqsubseteq w^{\prime}$ if and only if $w^{\prime}\sqsubseteq w$. We introduce the notation
$\Phi(\mathcal{W})\coloneqq\{\mathbf{O}_{\mathcal{W},w}:\operatorname{Im}(w)=\mathcal{W}\}$
and stress that this is a set of pairwise disjoint open subsets of $\Phi(X)$ by the remark above.
The following proposition shows that more is true: $\Phi(\mathcal{W})$ is a quasi-partition whose mesh is controlled by that of $\mathcal{W}$.
@@
Let $X$ be a Peano continuum, and $\mathcal{W}$ be a brick partition of $X$. Fix a compatible metric $d$ on $X$ and the induced metric $d_{\Phi}$ on $\Phi(X)$. Then $\Phi(\mathcal{W})$ is a quasi-partition of $\Phi(X)$ such that $\mathrm{mesh}_{d_{\Phi}}(\Phi(\mathcal{W}))\leq\mathrm{mesh}_{d}(\mathcal{W})$. Moreover, whenever $\mathcal{U}$ is a brick partition such that $\mathcal{U}\preceq\mathcal{W}$, then $\Phi(\mathcal{U})\preceq\Phi(\mathcal{W})$.
The proof of the proposition relies on the following lemma.
Lemma 5.3.
Let $X$ be a Peano continuum and $\mathcal{U}$ a brick partition of $X$. If $\mathcal{C},\mathcal{C}^{\prime}\in\Phi(X)$ are such that
(5.3) $\{\operatorname{Star}(K,\mathcal{U}):K\in\mathcal{C}\}\subseteq\{\operatorname{Star}(K^{\prime},\mathcal{U}):K^{\prime}\in\mathcal{C}^{\prime}\},$
then $d_{\Phi}(\mathcal{C},\mathcal{C}^{\prime})\leq\mathrm{mesh}_{d}(\mathcal{U})$.